DIRC Operations Weekly Report
Period: Wednesday 2007/3/04 - Wednesday 2007/4/10
Oncall: Alejandro Perez
Fast Monitoring: Alejandro Perez, Bradley Wogsland
PEP peak lumi: ~8.3 x 1033
(previous week: ~8.3 x 1033)
BaBar integrated lumi: 0.80 fb-1
(previous week: 1.97 fb-1)
Run 6 PEP delivery: 0.92 fb-1
(previous week: 2.07 fb-1)
Run 6 total delivery: ~24.99 fb-1
Run 6 total BaBar recorded: 23.97 fb-1
including 0 fb-1 off-peak data
PEP activities this week:
PEP-II down due to broken water pipe shut down BaBar's cooling water from 4/3 to 4/5
PEP-II down due to vacum leak from 4/7 to 4/10.
BaBar activities/issues this week:
- New plot for FM (Nicolas) about FX error 11.
- Identified one noisy channel sector 6, DFB 9, TDC 3, channel 3 that maybe have to maskedout.
- DFB#7 in sector #8 some times does not show up in the calibration plots since 4/6. The TDCs mean time for this DFB have been shifting. A standalone calibration on 4/9 showed that the timing distributions for all channels of this DFB have lower occupancy (a factor of 10 lower) compared to the other DFBs: still trying to undestand it.
logbook for details
- Replace two DFBs, one in sector 0 and one in sector 1 with a group of 8 innefficient or
flaky channels (require bridge).
- Replace new DFB in sector 8 with 18 (TDC chip 0) often dead channels and figure
out what is killing DFBs in that slot.
- Found a mini X-tree in sector 7 HV group 55 than only shows up during taking cosmics.
Jose, Alejandro and Bradley rotating backup & on-call commissioning